OpenStudy (anonymous):

!!!Solve!!!Ashley uses the rowing machine and the stair machine exercise program. Her trainer wants to do an exercise program that meets these 2 conditions 1) Ashley exercise for 40 minutes, dividing her time between the stair machine and the rowing machine. 2) Ashley will spend twice as much time on the stair machine as on the rowing machine. Miguel also uses the stair machine and the rowing machine at the gym for an exercise program. His trainer wants him on an exercise program that meets these 3 conditions. 1)Miguel will spend at most 60 minutes working out at the gym CONT.....

OpenStudy (anonymous):

2) Miguel will exercise for at least 30 minutes, dividing his time between the stair machine and the rowing machine. 3) Miguel will spend at least twice as much time on the stair machine as on the rowing machine Questions Part 1 a Define 2 variables to rep the amount of time ashley can spend on the two machines b Write a system of 2 equations that describe the relationships between the amount of time ashley can spend on the 2 machines. c. Solve the system of equations d Interpret the solution

Vocaloid (vocaloid):

let's start by defining our variables for ashley s = number of minutes on stair machine r = number of minutes on rowing machine using our two conditions, we have: s + r = 40 and s = 2r solve the system for s and r

OpenStudy (anonymous):

It would be 2r+r=40 then 3r=40, 40/3=answer